# Geo-Sign: Hyperbolic Contrastive Regularisation for Geometrically Aware Sign Language Translation

### v1

@@ 1 Introduction
Sign Languages are rich, multi-channel linguistic systems where meaning is conveyed through a composition of movements involving the upper body, hands, face, and mouth. Automatic Sign Language Translation (SLT) is an established research area focused on developing technologies to convert these visual expressions directly into text. While Sign Languages are expressed via fluid multi-articulator kinematics, a persistent challenge for SLT methods lies in creating feature representations that concurrently preserve fine-grained, local details (e.g., subtle finger configurations) while embedding the global structure inherent in larger, overarching body motions. Effectively modelling these multi-scale and relational dynamics within a suitable geometric embedding space remains a central hurdle.
@@
Spatio-Temporal Graph Convolutional Networks (ST-GCNs) offer a natural way to encode these hierarchical relationships by treating the body’s joints and bones as nodes and edges in a graph [74]. However, when their learned representations are projected into standard Euclidean geometry for processing via a Language Model (LM), essential fine-grained relational distances and movements can become blurred. For instance, the sign for “water” in American Sign Language (ASL) is communicated by forming a W shape with the fingers and tapping the chin twice (a fine-grained, "leaf-level" articulation), immediately followed by a sweeping hand movement away from the body (a "branch-level" gesture). When these features are aggregated in Euclidean geometry, the large translation and rotation of the wrist could dominate the vector’s norm, effectively “pulling” the embedding toward the global motion and compressing the subtle finger tap into a vanishing tail. Consequently, two signs that differ only in the timing or precision of that tap, which may be critical to lexical meaning, can become nearly indistinguishable once projected into flat Euclidean space.
@@
Large vision-based models [23, 22, 69, 21] appear to be able to implicitly learn these hierarchical structures through extensive video pre-training and visual inductive biases. However, they do so at significant computational cost and with privacy concerns, as they retain identifiable facial and background details that skeletal representations inherently discard.
@@
Figure 1: Geo-Sign’s hyperbolic framework: (Left) Skeletal features from ST-GCN’s for different body parts are projected into a Poincaré ball whose curvature is learned, while the original branch fuses the features for processing via the MT5 language model. (Pooled) The pose features are aggregated via Frechet Mean in Eq.1, while the text embeddings from the final layer of the MT5 model are pooled and projected to the hyperbolic manifold. Geodesic distance between the text embedding and the mean pose features are minimised for positive samples using the contrastive loss in Eq.5. (Token) Alternatively, hyperbolic pose features are used as attention queries against all text tokens to minimize the distance of the text to all positive pose samples. (Right Lower) A representation of the Poincaré disk demonstrating the difference between Token, and Pooled methods in the tangent space.
@@
This work introduces hyperbolic geometry as a means to fundamentally enhance skeletal representations for SLT. Unlike Euclidean space, where volume grows polynomially with radius and can flatten hierarchical structures, hyperbolic manifolds exhibit exponential volume growth. This property is naturally suited to encoding the compositional, tree-like structures found in sign language kinematics. As illustrated in Figure 1, in the Poincaré ball model $\mathbb{B}_{c}^{d_{\text{hyp}}}$ (with curvature $\kappa=-c<0$), distances between points near the boundary expand exponentially relative to their Euclidean separation. This provides ample "space" to distinguish nuanced motions (e.g., an open versus a closed fist), while regions near the origin behave more like Euclidean space, suitable for representing broader phrase-level semantics. A key aspect of our approach is that we learn the curvature parameter $c$ end-to-end via Riemannian optimization. This allows the manifold to dynamically adapt its "zoom level": a more negative curvature $\kappa$ (larger $c$) amplifies the separation of fine-grained motions, whereas a milder curvature helps preserve sentence-level coherence.
@@
Experiments on the CSL-Daily benchmark [80] demonstrate Geo-Sign’s efficacy. Our skeletal-based approach not only achieves a +1.81 BLEU4 and +3.03 ROUGE score over state-of-the-art pose-based methods but also matches the performance of comparable vision-based networks. We also present the first method to surpass SOTA gloss based methods (with respect to the ROUGE score) with a gloss-free approach, highlighting the potential of geometrically-aware representations.
@@ 2.1 Sign Language Translation (SLT)
Early SLT methods often involved a two-stage process: recognizing sign glosses (individual lexical units of sign language grammar) and then translating the gloss sequence into the target language [6, 82, 24, 25, 64, 65, 45, 46, 46, 66, 67]. However, this intermediate representation can lead to information loss, while gloss transcriptions are limited in availability. Consequently, end-to-end sequence-to-sequence models have become the dominant paradigm [5]. Initial approaches utilized Recurrent Neural Networks (RNNs) like LSTMs or GRUs, often with attention mechanisms [58, 20]. More recently, Transformer architectures [6] have demonstrated superior performance in capturing long-range dependencies and context [84, 57, 30], enabling direct video-to-text translation [69, 21, 15, 63, 10]. Many recent state-of-the-art architectures leverage large pre-trained language models, such as T5 variants, fine-tuned for the task of SLT [9, 82]. These often rely on large pre-trained visual encoders, with incremental improvements seen by upgrading visual backbones from ResNet [81], to I3D [62], and more recently to ViT variants like DINO [69]. However, as these backbones increase in size, they can limit the number of frames processed concurrently due to quadratically scaling resource demands.
@@ 2.2 Skeletal Representations for Sign Language and Action Recognition
Graph Convolutional Networks (GCNs) and particularly Spatio-Temporal GCNs (ST-GCNs) have shown great promise by explicitly modelling the spatial structure of the skeleton and its temporal dynamics [74, 71, 14, 72, 54]. However, the quality of skeletal data is heavily dependent on the accuracy of the underlying pose estimation algorithms [29]. Furthermore, skeletal data might discard subtle visual cues present in RGB video that could be important for disambiguation. While multi-modal fusion (RGB + pose) has been explored to combine the strengths of both modalities [50, 59, 83], it typically increases computational cost. Our work focuses on enhancing the representational power of skeletal data itself by embedding it in hyperbolic space, aiming to improve its discriminability for SLT without resorting to RGB fusion.
@@ 3.2.2 Part-Specific Projection to Poincaré Ball
3.3 Hyperbolic Contrastive Alignment Strategies
@@ 3.3.1 Strategy 1: Global Semantic Alignment (Pooled Method)
5:     for $p=1$ to $N$ do $\mathbf{v}_{\text{agg}}\leftarrow\mathbf{v}_{\text{agg}}+w_{p}\log_{\bm{\mu}^{(k)}}^{c}(\mathbf{h}_{p})$.
6:     end $\triangleright$ Sum weighted log-mapped vectors
7:     $\bm{\mu}^{(k+1)}\leftarrow\exp_{\bm{\mu}^{(k)}}^{c}(\mathbf{v}_{\text{agg}})$. $\triangleright$ Update mean via exponential map
8:     Project $\bm{\mu}^{(k+1)}$ into $\mathbb{B}_{c}^{d_{\text{hyp}}}$ if numerically necessary.
9:     if $d_{\mathbb{B}_{c}}(\bm{\mu}^{(k+1)},\bm{\mu}^{(k)})<\epsilon_{\text{tol}}$ then break.
10:     end $\triangleright$ Check convergence
11:end for
12:Estimated Fréchet mean $\bm{\mu}_{\text{pose}}=\bm{\mu}^{(k+1)}$.
3.3.2 Strategy 2: Fine-Grained Part-Text Alignment via Hyperbolic Attention (Token Method)
This strategy aligns individual pose parts (queries $\{\mathbf{h}_{p}\}$) with contextually relevant text segments, enabling more detailed semantic grounding.
  • Pose Embeddings $\{\mathbf{p}_{k}\}$: The set of $K$ individual hyperbolic part embeddings $\{\mathbf{h}_{p}\}$ from Eq. 4.
  • Text Embeddings $\{\mathbf{t}_{k}\}$: For each query $\mathbf{h}_{p}$, a contextual text embedding $\mathbf{c}_{p}\in\mathbb{B}_{c}^{d_{\text{hyp}}}$ is generated via hyperbolic attention:
    1. (a)
      Hyperbolic Tokenization: Euclidean text token embeddings are projected to $\mathbb{B}_{c}^{d_{\text{hyp}}}$ (similar to Eq. 4) yielding a sequence of hyperbolic token embeddings.
    2. (b)
      Hyperbolic Attention: These are then transformed using a learnable Möbius matrix-vector product, $\mathbf{M}\otimes_{c}\mathbf{v}=\exp_{\mathbf{0}}^{c}(\mathbf{M}\log_{\mathbf{0}}^{c}(\mathbf{v}))$ (where $\log_{\mathbf{0}}^{c}$ is inverse of Eq. 3), followed by a learnable Möbius addition (Eq. 2) with a bias vector. These are hyperbolic analogues of affine transformations. Attention scores are $-d_{\mathbb{B}_{c}}(\mathbf{h}_{p},\text{transformed key})$. Softmax (masked for padding) yields attention weights. $\mathbf{c}_{p}$ is then computed as the hyperbolic weighted midpoint $\mu$ of the original (untransformed) hyperbolic token values using these weights, providing a geometrically sound aggregation.
The final $\mathcal{L}_{\text{hyp\_reg}}$ for this strategy is the average of $K$ individual contrastive losses (Eq. 5), one for each $(\mathbf{h}_{p},\mathbf{c}_{p})$ pair.
@@ 4 Experiments
We evaluate Geo-Sign on the CSL-Daily dataset [82, 80], a large-scale corpus for Chinese Sign Language to Chinese text translation, comprising over 20,000 videos. Translation quality is assessed using BLEU [51] (B-1, B-4) and ROUGE-L [39] (R-L) scores where a higher percentage represents a more accurate translation.
@@ 4.1 Experimental Setup
Our framework builds upon the Uni-Sign architecture [38], using its pre-trained ST-GCN weights (trained on skeletal features from the CSL-News dataset) and an mT5 model [73] as the language decoder. Following Uni-Sign’s fine-tuning protocol, which involves 40 epochs of supervised finetuning on CSL-Daily with fused skeletal and RGB features, we remove the RGB encoder and instead apply our hyperbolic regularisation. This allows for a fair comparison of the impact of our geometric regularisation. We investigate both the "Pooled Method" (Strategy 1) and the "Token Method" (Strategy 2) for hyperbolic alignment. To assess the specific contribution of hyperbolic geometry, we also compare against a "Euclidean regularisation" baseline, where the contrastive loss operates on Euclidean projections to the Poincaré ball where curvature is minimal (0.001) and approximately Euclidean. Key hyperparameters for the hyperbolic components (initial curvature $c=1.5$, dimension $d_{\text{hyp}}=256$, and $\alpha=0.70$) are minimally tuned on the development set (further details in the appendix).
4.2 Quantitative Results and Ablation Studies
Section 4.2 presents our main results on the CSL-Daily test set, comparing Geo-Sign with prior art and baselines. Our Geo-Sign (Hyperbolic Token) model, using only pose data, achieves a test BLEU-4 of 27.42% and ROUGE-L of 57.95%. This represents a significant improvement of +1.81 BLEU-4 and +3.03 ROUGE-L over the strong Uni-Sign (Pose) baseline (25.61% BLEU-4, 54.92% ROUGE-L). Notably, this performance surpasses all other reported gloss-free pose-only methods and is competitive with, or exceeds, several RGB-only and even some gloss-based methods, underscoring the efficacy of our geometric regularisation. The Geo-Sign (Hyperbolic Pooled) variant also outperforms the Euclidean regularisation methods and the Uni-Sign pose baseline, demonstrating the general benefit of hyperbolic geometry. The "Euclidean Token" regularisation already shows improvement over the Uni-Sign baseline, suggesting the contrastive alignment itself is beneficial, but the further gains from hyperbolic geometry are substantial.
@@ 4.2 Quantitative Results and Ablation Studies
Ablation studies on the CSL-Daily test set for our best performing Geo-Sign (Hyperbolic Token) model are presented in Table 4. We investigate the impact of the initial hyperbolic curvature $c$ and the loss blending factor $\alpha$. For curvature $c$ (with $\alpha=0.7$), setting $c=0.001$ effectively makes the projection Euclidean (as $\tanh(x)\approx x$ for small $x$, which means almost zero hyperbolic warping). We observe that increasing curvature from this Euclidean-like baseline ($c=0.001$, BLEU-4 25.91%) generally improves performance, with optimal BLEU-4 (27.42%) achieved at $c=1.5$. ROUGE-L peaks at $c=2.0$ (58.08%), though BLEU-4 slightly dips to 27.25%, suggesting a trade-off. This indicates that a significant degree of negative curvature is beneficial for capturing sign language structure. For the loss blending factor $\alpha$ (with $c=1.5$), a value of $\alpha=0.7$ (i.e., 30% weight to the hyperbolic loss) yields the best BLEU-4 (27.42%) and ROUGE-L (57.95%). Lower or higher $\alpha$ values result in decreased performance, indicating that the hyperbolic regularisation provides a substantial complementary signal to the primary translation loss, but should not entirely dominate it during the 40 epochs of fine-tuning.
Table 2: Ablation experiments for Geo-Sign (Hyperbolic Token) on the CSL-Daily test set, varying initial curvature $c$ (with $\alpha=0.7$) and loss blending factor $\alpha$ (with $c=1.5$).
{subcaptionblock}
0.48 Curvature ($c$) BLEU-4 ROUGE-L 0.00 (Euclidean) 25.98 53.93 0.10 26.56 57.56 0.50 26.34 56.30 1.00 27.04 57.67 1.50 27.42 57.95 2.00 27.25 58.08 {subcaptionblock}0.48 $\alpha$ BLEU-4 ROUGE-L 0.10 25.74 56.20 0.50 26.79 57.38 0.70 27.42 57.95 0.90 26.92 57.67
Table 3: Impact of Initial Curvature $c$.
Table 4: Impact of Loss Blending $\alpha$.

### v2

@@ 1 Introduction
Sign Languages are rich, multi-channel linguistic systems where meaning is conveyed through a composition of movements involving the upper body, hands, face, and mouth. Automatic Sign Language Translation (SLT) is an established research area focused on developing methods to convert these visual expressions directly into text. While Sign Languages are expressed via fluid multi-articulator kinematics, a persistent challenge for SLT methods lies in creating feature representations that concurrently preserve fine-grained, local details (e.g., subtle finger configurations) while embedding the global structure inherent in larger, overarching body motions. Effectively modelling these multi-scale and relational dynamics within a suitable geometric embedding space remains a central hurdle.
@@
Spatio-Temporal Graph Convolutional Networks (ST-GCNs) offer a natural way to encode these hierarchical relationships by treating the body’s joints and bones as nodes and edges in a graph [79]. However, when their learned representations are projected into standard Euclidean geometry for processing via a Large Language Model (LLM), essential fine-grained relational distances and movements can become blurred. For instance, the sign for “water” in American Sign Language (ASL) is communicated by forming a W shape with the fingers and tapping the chin twice (a fine-grained, "leaf-level" articulation), immediately followed by a sweeping hand movement away from the body (a "branch-level" gesture). When these features are aggregated in Euclidean geometry, the large translation and rotation of the wrist could dominate the vector’s norm, effectively “pulling” the embedding toward the global motion and compressing the subtle finger tap into a vanishing tail. Consequently, two signs that differ only in the timing or precision of that tap, which may be critical to lexical meaning, can become nearly indistinguishable once projected into flat Euclidean space.
@@
Large vision-based models [23, 26, 22] appear to be able to implicitly learn these hierarchical structures through extensive video pre-training and visual inductive biases. However, they do so at significant computational cost and with privacy concerns, as they retain identifiable facial and background details that skeletal representations inherently discard.
@@
Figure 1: Geo-Sign’s hyperbolic framework: (Left) Skeletal features from ST-GCN’s for different body parts are projected into a Poincaré ball whose curvature is learned, while the original branch fuses the features for processing via the MT5 language model. (Pooled) The pose features are aggregated via Frechet Mean in Eq.1, while the text embeddings from the final layer of the MT5 model are pooled and projected to the hyperbolic manifold. Geodesic distance between the text embedding and the mean pose features are minimised for positive samples using the contrastive loss in Eq.5. (Token) Alternatively, hyperbolic pose features are used as attention queries against all text embeddings to generate a pose-contextual text embedding. Note the movement of the text features $c_{pi}$ in grey towards the pose feature in blue. (Right) A representation of the Poincaré disk demonstrating the difference between Token, and Pooled methods in the tangent space.
@@
This work introduces hyperbolic geometry as a means to fundamentally enhance skeletal representations for SLT. Unlike Euclidean space, where volume grows polynomially with radius and can flatten hierarchical structures, hyperbolic manifolds exhibit exponential volume growth. This property is naturally suited to encoding the compositional, tree-like structures found in sign language kinematics. As illustrated in Figure˜1, in the Poincaré ball model $\mathbb{B}_{c}^{d_{\text{hyp}}}$ (with curvature $\kappa=-c<0$), distances between points near the boundary expand exponentially relative to their Euclidean separation. This provides ample "space" to distinguish nuanced motions (e.g., an open versus a closed fist), while regions near the origin behave more like Euclidean space, suitable for representing broader phrase-level semantics. A key aspect of our approach is that we learn the curvature parameter $c$ end-to-end via Riemannian optimization. This allows the manifold to dynamically adapt its "zoom level": a more negative curvature $\kappa$ (larger $c$) amplifies the separation of fine-grained motions, whereas smoother curvature helps preserve sentence-level coherence.
@@
Experiments on the CSL-Daily benchmark [88] demonstrate Geo-Sign’s efficacy. Our skeletal-based approach not only achieves a +1.81 BLEU4 and +3.03 ROUGE score over state-of-the-art pose-based methods but also matches the performance of comparable vision-based networks. We demonstrate that our method extends to American Sign Language and Isolated Sign Language Recognition. Finally, we also present the first method to surpass SOTA gloss based methods (with respect to the ROUGE score) with a gloss-free approach, highlighting the potential of geometrically-aware representations.
@@ 2.1 Sign Language Translation (SLT)
Early SLT methods often involved a two-stage process: recognizing sign glosses (individual lexical units of sign language grammar) and then translating the gloss sequence into the target language [6, 89, 25, 27, 69, 70, 50, 49, 49, 72, 71]. However, this intermediate representation can lead to information loss, while gloss transcriptions are limited in availability. Consequently, end-to-end sequence-to-sequence models have become the dominant paradigm [5]. Initial approaches utilized Recurrent Neural Networks (RNNs) like LSTMs or GRUs, often with attention mechanisms [63, 21]. More recently, Transformer architectures [6] have demonstrated superior performance in capturing long-range dependencies and context [91, 62, 32], enabling direct video-to-text translation [74, 22, 15, 68, 9]. Many recent state-of-the-art architectures leverage large pre-trained language models, such as T5 variants, fine-tuned for the task of SLT [11, 89]. These often rely on large pre-trained visual encoders, with incremental improvements seen by upgrading visual backbones from ResNet [87], to I3D [67], and more recently to ViT variants like DINO [74, 75]. However, as these backbones increase in size, they can limit the number of frames processed concurrently due to quadratically scaling resource demands.
@@ 2.2 Skeletal Representations for Sign Language and Action Recognition
Graph Convolutional Networks (GCNs) and particularly Spatio-Temporal GCNs (ST-GCNs) have shown great promise by explicitly modelling the spatial structure of the skeleton and its temporal dynamics [79, 76, 14, 77, 58]. However, the quality of skeletal data is heavily dependent on the accuracy of the underlying pose estimation algorithms [30]. Furthermore, skeletal data might discard subtle visual cues present in RGB video that could be important for disambiguation. While multi-modal fusion (RGB + pose) has been explored to combine the strengths of both modalities [54, 64, 90, 75], it typically increases computational cost. Our work focuses on enhancing the representational power of skeletal data itself by embedding it in hyperbolic space, aiming to improve its discriminability for SLT without resorting to RGB fusion.
@@ 3.2.2 Part-Specific Projection to Poincaré Ball
3.3 Hyperbolic Contrastive Loss
@@
3.4 Alignment Methods
We present two methods for selecting features for alignment which offer benefits and trade-offs. The first takes the geometric mean of the pose and the text embeddings which comes at greater computational efficiency but decreased accuracy. The second method uses poses as individual queries over the text embeddings in hyperbolic space. We then compute the distance between each pose and modified token pair. This improves translation accuracy but incurs additional memory and inference costs (Full details provided in the appendix.)
@@ 3.4.1 Strategy 1: Global Semantic Alignment (Pooled Method)
5:  for $p=1$ to $N$ do
6:   $\mathbf{v}_{\text{agg}}\leftarrow\mathbf{v}_{\text{agg}}+w_{p}\log_{\bm{\mu}^{(k)}}^{c}(\mathbf{h}_{p})$. $\triangleright$ Sum weighted log-mapped vectors
7:  end for
8:  $\bm{\mu}^{(k+1)}\leftarrow\exp_{\bm{\mu}^{(k)}}^{c}(\mathbf{v}_{\text{agg}})$. $\triangleright$ Update mean via exponential map
9:  Project $\bm{\mu}^{(k+1)}$ into $\mathbb{B}_{c}^{d_{\text{hyp}}}$ if numerically necessary.
10:  if $d_{\mathbb{B}_{c}}(\bm{\mu}^{(k+1)},\bm{\mu}^{(k)})<\epsilon_{\text{tol}}$ then $\triangleright$ Check convergence
11:   break $\triangleright$ Exit loop on convergence
12:  end if
13:end for
[…]
15:Estimated Fréchet mean $\bm{\mu}_{\text{pose}}$.
3.4.2 Strategy 2: Fine-Grained Part-Text Alignment (Token Method)
This strategy aligns individual pose parts (Queries $\bm{h}_{p}$) with relevant text segments via hyperbolic attention. For each $\bm{h}_{p}\in\mathbb{B}_{c}^{d_{\text{hyp}}}$ (from Eq. 4), a specific context vector $\bm{c}_{p}\in\mathbb{B}_{c}^{d_{\text{hyp}}}$ is computed.
  • Tokenization: First, the sequence of $T$ Euclidean text token embeddings, $\{\bm{e}_{t}\}_{t=1}^{T}$, is projected into the Poincaré ball. This yields a sequence of hyperbolic embeddings $V=\{\bm{v}_{t}\}_{t=1}^{T}$, which serve as the values and hold the original semantic meaning of each token.
  • Key Transformation: The value sequence $V$ is transformed by a learnable Möbius affine transformation (using $\bm{M}$ and $\bm{b}$) to create a sequence of keys $K=\{\bm{k}_{t}\}_{t=1}^{T}$:
    $\bm{k}_{t}=(\bm{M}\otimes\bm{v}_{t})\oplus\bm{b}$ (6)
  • Attention Scores: Scores $s_{pt}$ are computed as the negative geodesic distance between each pose query $\bm{h}_{p}$ and each text key $\bm{k}_{t}$. A smaller distance signifies greater relevance.
    $s_{pt}=-d_{\mathbb{B}_{c}}(\bm{h}_{p},\bm{k}_{t})$ (7)
  • Context Vector: These scores are normalized via softmax (with masking for padding) to produce attention weights $\alpha_{pt}$. The final context vector $\bm{c}_{p}$ is the hyperbolic weighted midpoint ($\mu$) of the original values $V$ weighted by the pose-shifted attention embeddings:
    $\bm{c}_{p}=\mu_{\mathcal{B}_{c}}\left(\{\bm{v}_{t}\}_{t=1}^{T},\{\alpha_{pt}\}_{t=1}^{T}\right)$ (8)
    The final $\mathcal{L}_{\text{hyp\_reg}}$ is the average of $K$ individual contrastive losses (Eq. 5), one for each $(\bm{h}_{p},\bm{c}_{p})$ alignment pair.
@@ 4 Experiments
We evaluate Geo-Sign on Chinese Sign Language (CSL) and American Sign Language (ASL). For CSL we use the CSL-Daily dataset [89, 88], a large-scale corpus for Chinese Sign Language to Chinese text translation, comprising over 20,000 videos. For American Sign Language we perform translation experiments on How2Sign [15] and isolated sign language recognition experiments on WLASL2000 [39].
Translation quality is assessed using BLEU [55] (B-1, B-4) and ROUGE-L [42] (R-L) scores where a higher percentage represents a more accurate translation.
@@ 4.1 Experimental Setup
Our framework builds upon the Uni-Sign architecture [41], using its pre-trained ST-GCN weights (trained on skeletal features from the CSL-News dataset [41] for CSL and the YTASL [68] dataset for ASL) and an mT5 model [78] as the language decoder. Following Uni-Sign’s fine-tuning protocol, which involves 40 epochs of supervised finetuning on CSL-Daily or YTASL, with fused skeletal and RGB features, we remove the RGB encoder and instead apply our hyperbolic regularisation. This allows for a fair comparison of the impact of our geometric regularisation. We investigate both the "Pooled Method" (Strategy 1) and the "Token Method" (Strategy 2) for hyperbolic alignment. To assess the specific contribution of hyperbolic geometry, we also compare against a "Euclidean regularisation" baseline, where the contrastive loss operates on Euclidean projections to the Poincaré ball where curvature is minimal (0.001) and approximately Euclidean. Key hyperparameters for the hyperbolic components (initial curvature $c=1.5$, dimension $d_{\text{hyp}}=256$, and $\alpha=0.70$) are minimally tuned on the development set (further details in the appendix).
4.2 Results on Chinese Sign Language (CSL)
Section˜4.2.1 presents our main results on the CSL-Daily test set, comparing Geo-Sign with prior art and baselines. Our Geo-Sign (Hyperbolic Token) model, using only pose data, achieves a test BLEU-4 of 27.42% and ROUGE-L of 57.95%. This represents a significant improvement of +1.81 BLEU-4 and +3.03 ROUGE-L over the strong Uni-Sign (Pose) baseline (25.61% BLEU-4, 54.92% ROUGE-L). Notably, this performance surpasses all other reported gloss-free pose-only methods and is competitive with, or exceeds, several RGB-only and even some gloss-based methods, underscoring the efficacy of our geometric regularisation. The Geo-Sign (Hyperbolic Pooled) variant also outperforms the Euclidean regularisation methods and the Uni-Sign pose baseline, demonstrating the general benefit of hyperbolic geometry. The "Euclidean Token" regularisation already shows improvement over the Uni-Sign baseline, suggesting the contrastive alignment itself is beneficial, but the further gains from hyperbolic geometry are substantial.
4.2.1 Results on American Sign Language
In Table 4.2.1 we show results on Sign Language Translation for American Sign Language on the How2Sign [15] dataset. Our method shows increased performance over all pose based methods but performs marginally worse than the best RGB method [60] which benefits from a longer pre-training duration and scale. In Table 4.2.1 we also compare our approach on Isolated Sign Language Recognition (ISLR) with the WLASL2000 [39] dataset. For isolated recognition our method shows a small improvement in Top-1 Accuracy for both instance (+0.12) and class-level (+0.57).
@@ 4.2.1 Results on American Sign Language
Table 2: Sign Language Translation (SLT) results on the How2Sign dataset. Metrics are BLEU (B-1, B-4) and ROUGE-L (R-L). Higher is better.
Method Modality Test Set
Pose RGB B-1 B-4 R-L
[…]
GloFE-VN [43] 14.9 2.2 12.6
YouTube-ASL [68] 37.8 12.4 -
MSLU [90] 20.1 2.4 17.2
SLT-IV [67] 34.0 8.0 -
C2RL [9] 29.1 9.4 27.0
FLa-LLM [10] 29.8 9.7 27.8
SignMusketeers [20] 41.5 14.3 -
SSVP-SLT [60] 43.2 15.5 38.4
[…]
Uni-Sign (Pose) 40.4 14.5 34.3
Uni-Sign (Pose+RGB) 40.2 14.9 36.0
Geo-Sign (Token) 40.8 15.1 35.4
Table 3: Isolated Sign Language Recognition (ISLR) results on WLASL2000. We report Top-1 Accuracy for Per-Instance (P-I) and Per-Class (P-C). from [23].
Method Mod. Test Acc. (%)
P RGB P-I P-C
[…]
ST-GCN [80] 34.40 32.53
SignBERT [23] 39.40 36.74
HMA [24] 37.91 35.90
BEST [86] 46.25 43.52
SignBERT+ [26] 48.85 46.37
MSLU [90] 56.29 53.29
NLA-SLR [93] 61.05 58.05
Sign-Rep [75] 61.05 58.89
[…]
Uni-Sign (Pose) 63.13 60.90
Uni-Sign (Pose+RGB) 63.52 61.32
Geo-Sign (Token) 63.64 61.89
4.2.2 Ablation Studies
Ablation studies on the CSL-Daily test set for our best performing Geo-Sign (Hyperbolic Token) model are presented in Table˜4. We investigate the impact of the initial hyperbolic curvature $c$ and the loss blending factor $\alpha$. For curvature $c$ (with $\alpha=0.7$), setting $c=0.001$ effectively makes the projection Euclidean (as $\tanh(x)\approx x$ for small $x$, which means almost zero hyperbolic warping). We observe that increasing curvature from this Euclidean-like baseline ($c=0.001$, BLEU-4 25.91%) generally improves performance, with optimal BLEU-4 (27.42%) achieved at $c=1.5$. ROUGE-L peaks at $c=2.0$ (58.08%), though BLEU-4 slightly dips to 27.25%, suggesting a trade-off. This indicates that a significant degree of negative curvature is beneficial for capturing sign language structure. For the loss blending factor $\alpha$ (with $c=1.5$), a value of $\alpha=0.7$ (i.e., 30% weight to the hyperbolic loss) yields the best BLEU-4 (27.42%) and ROUGE-L (57.95%). Lower or higher $\alpha$ values result in decreased performance, indicating that the hyperbolic regularisation provides a substantial complementary signal to the primary translation loss, but should not entirely dominate it during the 40 epochs of fine-tuning. Finally, we assess the robustness of our approach to pose perturbation and poor pose estimation. To do so, we add Gaussian noise to the pose embeddings before embedding via the ST-GCN. Our method shows how hyperbolic regularisation improves robustness to pose noise compared to the Uni-Sign Euclidean baseline. We attribute this to the larger geodesic margins between pose embeddings in the hyperbolic space, making the model less susceptible to noisy perturbation.
Table 4: Ablation studies for Geo-Sign on the CSL-Daily test set, examining (a) initial curvature $c$, (b) loss blending factor $\alpha$, and (c) robustness of poses to Gaussian noise. We demonstrate that hyperbolic regularisation improves robustness to pertubation and poor pose estimation.
Curvature ($c$) B-4 R-L
0.00 (Euclidean) 25.98 53.93
0.10 26.56 57.56
0.50 26.34 56.30
1.00 27.04 57.67
1.50 27.42 57.95
2.00 27.25 58.08
(a) Impact of Curvature $c$.
$\alpha$ B-4 R-L
0.10 25.74 56.20
0.50 26.79 57.38
0.70 27.42 57.95
0.90 26.92 57.67
(b) Impact of $\alpha$.
Noise Geo-Sign Uni-Sign
0.00 27.42 26.25
0.01 26.30 (-4%) 24.14 (-8%)
0.02 24.60 (-10%) 21.50 (-18%)
0.03 19.07 (-30%) 14.40 (-45%)
0.04 11.63 (-58%) 7.20 (-73%)
0.05 5.98 (-78%) 3.01 (-89%)
(c) Noise Robustness (B-4).